Find dy/dx if y=1/(sinx cosx) differentiate this.>50 characters

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
y = 1 / sin x cosx
 
y = 2 / sin 2x
y = 2 cosec 2x
 
now
 
dy/dx = 2*2 ln(sin x))
 
 
Regards
Arun (askIITians foum expert)